The process and answer for the equation
10y+8y=15

Answer by tutorcecilia(2152)

10y+8y=15 [Combine like-terms]
18y=15 [Solve for y]
18y/18=15/18
y=15/18 [Reduce the fraction]
y=5/6
.
Check by plugging (y=5/6) back into the original equation
10y+8y=15
10(5/6)+8(5/6)=15
(50/6)+(40/6)=15
15=15 [Checks out]